HIGH RESOLUTION LASER SPECTROSCOPY OF Mg^{12}C^{12}CD, Mg^{13}C^{13}CH AND Mg^{12}C_{4}H

\underline{D. FORTHOMME}, C. LINTON, D. W. TOKARYK, Centre for Laser, Atomic, and Molecular Sciences and Physics Department, 8 Bailey Dr., University of New Brunswick, P.O. Box 4400, Fredericton, NB, Canada E3B 5A3; A .G. ADAM, A. D. GRANGER, Centre for Laser, Atomic, and Molecular Sciences and Chemistry Department, 30 Dineen Dr., University of New Brunswick, P.O. Box 4400, Fredericton, NB, Canada E3B 5A3.

Carbon and magnesium are abundant elements in the interstellar medium, so it is possible that carbon chain molecules containing a magnesium atom may exist in this environment^{1}.With this in mind, radical molecules of the form MgC_{2n}H (n = 1,2,3) have been frequent subjects of both experimental and theoretical studies^{a}^{b}^{c}^{d}^{e}^{f}. In this presentation we will discuss our high-resolution experiments of the Ã^{2}Π−X̃^{2}Σ^{+} transitions in the isotopologues Mg^{12}C^{12}CD and Mg^{13}C^{13}CH, which complement our earlier investigation of this spectrum in Mg^{12}C^{12}CH^{b}. The data permit us to determine the lengths of individual bonds to high precision. In addition, we have expanded on previous studies of the Ã^{2}Π − X̃^{2}Σ^{+} transition of Mg^{12}C_{4}H, conducted at medium resolution. The parameters obtained from our high-resolution spectra are compared with those obtained from theoretical structure calculations.

aH. Ding, C. Apetrei, L. Chacaga, J. P. Maier, Astrophys. J. 677 (2008) 348-352

bD. W. Tokaryk, A. G. Adam, W. S. Hopkins, J. Mol. Spectrosc. 230 (2005) 54-61

cD. E. Woon, Chem. Phys. Lett. 274 (1997) 299-305

dC. A. Thompson and L. Andrews, J. Am. Chem. Soc. 118 (1996) 10242-10249

eX. Guo, J. Zhang, J. Li, L. Jiang, J. Zhang, Chem. Phys 360 (2009) 27-31

fE. Chasovskikh, E. B. Jochnowitz, J. P. Maier, J. Phys. Chem. A. 112 (2008) 8686-8689